OpenStudy (anonymous):

HELP! Simplify the expression: 7sqrt7- 2sqrt28 A. 3sqrt28 B. -sqrt7 C. 5sqrt28 D. 3sqrt7

OpenStudy (johnweldon1993):

Hint: write it instead as... \[\large 7\sqrt{7} - 2\sqrt{7 \times 4}\] any better?

OpenStudy (anonymous):

i got d 3 squard 7

OpenStudy (anonymous):

D it is

OpenStudy (johnweldon1993):

So from that^ we can see that the square root of 4...is 2... Since we can simplify that √4 = 2...we now have \[\large 7\sqrt{7} - 2 \times 2\sqrt{7}\] \[\large 7\sqrt{7} - 4\sqrt{7}\] Now we just combine the two, \[\large 3\sqrt{7}\]

OpenStudy (anonymous):

Thank you so much! @johnweldon1993

OpenStudy (johnweldon1993):

No problem :)

OpenStudy (anonymous):

wow